### v1

@@ Abstract
Symmetry-informed machine learning can exhibit advantages over machine learning which fails to account for symmetry. Additionally, recent attention has been given to continuous symmetry discovery using vector fields which serve as infinitesimal generators for Lie group symmetries. In this paper, we extend the notion of non-affine symmetry discovery to functions defined by neural networks. We further extend work in this area by introducing symmetry enforcement of smooth models using vector fields. Finally, we extend work on symmetry discovery using vector fields by providing both theoretical and experimental material on the restriction of the symmetry search space to infinitesimal isometries.
@@
Symmetry Discovery, Machine Learning, Symmetry Enforcement, Killing Vectors
{MSCcodes}
62-07, 68T01
@@ 1 Introduction
Symmetry in the context of machine learning can be exploited to improve performance and/or computational efficiency [16, 3, 7, 26]. Recently, the detection of symmetry in data has been accomplished by means of infinitesimal generators, which has allowed for the discovery of symmetries which are not affine [14, 23]. Herein, we seek to extend the work of symmetry discovery using infinitesimal generators, demonstrate a means of enforcing symmetry in machine learning models using them, and further explore the discovery of infinitesimal isometries.
@@
Our method of discovering symmetry using infinitesimal generators is summarized as follows. First, given a dataset, certain quantities of interest, hereafter referred to as machine learning functions, are identified, such as an underlying probability distribution which generates the dataset, a classification or regression function, or a dimensionality reduction function. Second, we estimate tangent vector fields which annihilate the machine learning functions. The estimated vector fields themselves are infinitesimal generators of continuous symmetries. Thus, estimating the vector fields provides an indirect estimate of the continuous symmetries.
We offer two distinct methods of enforcing symmetry. Given previously [23], the first method uses the estimated vector fields to estimate functions which, in addition to the machine learning functions from which the vector fields were estimated, are annihilated by the vector fields. The discovered functions define a feature space which is invariant with respect to the continuous symmetries implied by the vector field generators. On this constructed feature space, the outcome of any machine learning task is, by definition, invariant with respect to the symmetries generated by the estimated vector fields. The second method of symmetry enforcement seeks to regularize the learning of a machine learning model by enforcing the vector field annihilation condition directly in the training of the model.
We also extend previous work on the restriction of the space of transformations to isometries. Previous work on symmetry discovery frequently restricts the types of symmetries being sought: common restrictions are affine symmetries [28] or classical Lie group transformations [10]. Recent work on symmetry discovery using infinitesimal generators has demonstrated potential in discovering more general transformations [14], though there is evidence that attempting to search for more generic infinitesimal generators can incur undesirable computational costs [23]. To address the shortfalls of both approaches, we propose restricting the search space to geometrically meaningful symmetries such as isometries. In principle, restrictions can be made to other types of transformations, such as conformal transformations, though at present we consider only the restriction to isometries.
@@
This paper is organized as follows. In the next section, we discuss previous work in symmetry detection and enforcement. In section 3, we discuss vector fields, flows, and isometries, the mathematics that underlies our methods. In section 4, we describe our methods, and in section 5, we present our experimental results. We then offer concluding remarks and outline future work in the area.
@@ 2.1 Symmetry Discovery
Early work on symmetry detection in machine learning focused primarily on detecting symmetry in image and video data [21, 24], where symmetries described by straight-line and rotational transformations were discovered. Other work has made strides in symmetry discovery by restricting the types of symmetries being sought. In one case, detection was limited to compact Abelian Lie groups [5] and used for the purpose of learning disentangled representations. Another case uses meta-learning to discover any finite symmetry group [30]. Finite groups have also been used in symmetry discovery in representation learning [1]. In physics-based applications of machine learning, a method has been constructed which is able to discover any classical Lie group symmetry [10]. Symmetry discovery of shapes has also been explored [12].
@@
Other work has focused on detecting affine transformation symmetries and encoding the discovered symmetries automatically into a model architecture. Three such methods identify Lie algebra generators to describe the symmetries, as we do herein. For example, augerino [2] attempts to learn a distribution over augmentations, subsequently training a model with augmented data. The Lie algebra convolutional network [8], which generalizes Convolutional Neural Networks in the presence of affine symmetries, uses infinitesimal generators represented as vector fields to describe the symmetry. SymmetryGAN [9] has also been used to detect rotational symmetry [28].
@@
Another notable contribution to efforts to detect symmetries of data is LieGAN. LieGAN is a generative-adversarial network intended to return infinitesimal generators of the continuous symmetry group of a given dataset [28]. LieGAN has been shown to detect continuous affine symmetries, including transformations from the Lorentz group. It has also been shown to identify discrete symmetries such as rotations by a fixed angle.
While most continuous symmetry detection methods attempt to discover symmetries which are affine transformations, the representation of infinitesimal generators using vector fields has led to the discovery of continuous symmetries which are not affine [14, 23]. In one case, the domains of image data and partial differential equations are examined in particular [14]. In the case for which this paper is an extension of [23], tabular data is examined, and the infinitesimal generators are either isometries or else usually contain polynomial coefficients.
@@
Continuous symmetry detection is more difficult than discrete symmetry detection [30] since the condition $f\circ S=f$ must hold for all values of the continuous parameter of $S$. This is corroborated by the increasingly complex methods used to calculate even simple symmetries such as planar rotations [2, 8, 28]. Some methods introduce discretization, where multiple parameter values are chosen and evaluated. LieGAN does this by generating various transformations from the same infinitesimal generator [28]. Introducing discretization increases the complexity of continuous symmetry detection, though any reasonable symmetry detection method must establish whether a continuous symmetry a finite subgroup of a continuous group has been discovered. We believe a vector field approach addresses the issue of discretization. Our vector field-based method reduces the required model complexity of continuous symmetry detection while offering means to detect symmetries beyond affine transformations.
@@ 2.2 Symmetry Enforcement
Previous work on enforcing symmetry in machine learning models can be categorized. Some methods seek to enforce symmetry by augmenting the training dataset according to known symmetries [3]: this is common practice in image classification tasks, where augmentations include rotated copies of images in the training set. Augerino attempts to enforce symmetry using augmented data, though the symmetries are discovered from the data rather than given a priori. Another established method of enforcing symmetry is feature averaging, which is thought to be generally more effective than data augmentation [16].
A growing number of methods seek to use symmetry to construct invariant and equivariant models without the need for augmented training data, and two previously-mentioned sources do this using infinitesimal generators [8, 28]. Some previous work addresses specific cases: the special case of compact groups is studied [4], and the case of equivariant CNNs on Homegeneous spaces is studied [6]. Other work speaks to the universality of invariant architectures [17, 13, 29].
Another method which may be seen as a type of symmetry enforcement is the deployment of Physics-Informed Neural Networks (PINNs) [20]. Here, model training is regularized using differential constraints which represent the governing equations for a physical system. Our method of symmetry enforcement is, in many respects, mathematically equivalent, though the differential constrains obtained using infinitesimal generators do not generally have the interpretation of defining governing equations for a physical system.
@@ 3 Background
In this section, we will give an overview of vector fields and flows, as well as their connection to symmetry. We will also discuss Killing vectors and isometries. Throughout, we assume a dataset $\mathcal{D}=\{x_{1},\dots,x_{r}\}$ with $x_{i}\in\mathbb{R}^{n}$. Various machine learning functions of interest may exist for a given dataset, such as an underlying probability distribution or class probabilities. Generally speaking, the object of symmetry detection in data is to discover functions $S:\mathbb{R}^{n}\to\mathbb{R}^{n}$ that preserve a machine learning function of interest: that is, for a machine learning function $f$, $f\circ S=f$. We consider a continuous symmetry of a particular machine learning function to be a transformation $S$ which is continuously parameterized, such as a rotation in a plane by an angle $\theta$, with $\theta$ being the continuous parameter. We consider a discrete symmetry of a machine learning function to be a transformation which is not continuously parameterized, such as a planar rotation by a fixed angle, or a reflection of a point in the plane about a straight line. This paper deals exclusively with continuous symmetries.
@@ 3.1 Vector Fields and Flows
We now provide some background on vector fields and their associated flows (1-parameter transformations). We refer the reader to literature on the subject for additional information [15]. Suppose that $X$ is a smooth [^1] That is, a vector field with $\mathcal{C}^{\infty}$ coefficient functions. (tangent) vector field on $\mathbb{R}^{n}$:
@@
For example, for $n=2$, if $f(x,y)=xy$ and $X=y\partial_{x}$, then $X(f)=y^{2}$. $X$ is also a derivation on the set of smooth functions on $\mathbb{R}^{n}$: that is, for smooth functions $f_{1},f_{2}:\mathbb{R}^{n}\to\mathbb{R}$ and $a_{1},a_{2}\in\mathbb{R}$,
(3) $X(a_{1}f_{1}+a_{2}f_{2})=a_{1}X(f_{1})+a_{2}X(f_{2}),\qquad X(f_{1}f_{2})=X(f_{1})f_{2}+f_{1}X(f_{2}).$
These properties are satisfied by derivatives. A flow on $\mathbb{R}^{n}$ is a smooth function $\Psi:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ which satisfies
@@
for all $s,t\in\mathbb{R}$ and for all $p\in\mathbb{R}^{n}$. A flow is a 1-parameter group of transformations. An example of a flow $\Psi:\mathbb{R}\times\mathbb{R}^{2}\to\mathbb{R}^{2}$ is
@@
For a given flow $\Psi$, one may define a (unique) vector field $X$ as given in Equation 2, where each function $\alpha^{i}$ is defined as
(6) $\alpha^{i}=\left(\frac{\partial\Psi}{\partial t}\right)\bigg{|}_{t=0}.$
Such a vector field is called the infinitesimal generator of the flow $\Psi$. For example, the infinitesimal generator of the flow given in Equation 5 is $-y\partial_{x}+x\partial_{y}$.
@@
Suppose that a solution $\mathbf{x}(t)$ to Equation 7 exists for all $t\in\mathbb{R}$ and for all initial conditions $\mathbf{x}_{0}\in\mathbb{R}^{n}$. Then the function $\Psi:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ given by
(8) $\Psi(t,\mathbf{x}_{0})=\mathbf{x}(t)$
is a flow. The infinitesimal generator corresponding to $\Psi$ is $X$. For example, to calculate the flow of $-y\partial_{x}+x\partial_{y}$, we solve
@@ 3.2 Vector Fields and Symmetry
We can now connect vector fields and flows with symmetry. A smooth function $f:\mathbb{R}^{n}\to\mathbb{R}$ is said to be $X$-invariant if $X(f)=0$ identically for a smooth vector field $X$. The function $f$ is $\Psi$-invariant if, for all $t\in\mathbb{R}$, $f=f(\Psi(t,\cdot))$ for a flow $\Psi$. If $X$ is the infinitesimal generator of $\Psi$, $f$ is $\Psi$-invariant if and only if $f$ is $X$-invariant. If the function $f$ is a machine learning function for a given data set, our strategy is to identify vector fields $X_{i}$ for which $f$ is invariant. Each vector field $X_{i}$ for which $f$ is invariant is associated, explicitly or implicitly, with a flow $\Psi_{i}$, each of which is a 1-parameter subgroup, the collection of which generate the symmetry group. By identifying continuous symmetries by means of $X(f)=0$, continuous symmetry detection is made to be of similar complexity to that of discrete symmetry detection, since no continuous parameters are present.
The price of this indirect characterization of continuous symmetry is often the loss of the explicit identification of the symmetry group, leaving one with only the infinitesimal generators and the symmetry group merely implied. However, infinitesimal generators can be used to construct an invariant machine learning model as follows. Suppose that $f:\mathbb{R}^{n}\to\mathbb{R}$ is a (smooth) function of interest for a given data set, and suppose that $X$ is a (smooth) vector field for which $f$ is $X$-invariant. With $X$ fixed, we can consider whether there exist additional functions $\{h^{i}\}_{i=1}^{m}$ that are functionally independent of each other (meaning one function cannot be written as a function of the others), each of which are $X$-invariant. If such functions can be found, we can define a new feature space using the functions $\{h^{i}\}$. That is, we can transform each data point in $\mathcal{D}$ to the point $(h^{1}(x_{i}),h^{2}(x_{i}),\dots,h^{m}(x_{i}))$ for each point $x_{i}$ expressed in the original coordinates $\{x^{i}\}$. By definition, the transformed data is invariant with respect to the symmetry group, whether that symmetry group is given explicitly or not.
The flow parameter can often be obtained explicitly as well, written as a function of the original coordinates. It is found by solving $X(\theta)=1$. Where the flow parameters of vector fields can be estimated, our methodology thus inspires symmetry-based feature engineering, with the new features consisting of flow parameters and invariant features.
For example, consider a data set consisting of points $\mathcal{D}=\{x_{i}\}_{i=1}^{r}$ in $\mathbb{R}^{3}$, where the features of each data point are given in Cartesian coordinates $(x,y,z)$. Now suppose that each data point $x_{i}$ satisfies the equation $f=0$, where $f=x^{2}+y^{2}-z$. The function $f$ can be considered to be a machine learning function for this data set, since each point in the data set lies on the level set $f=0$. We note that $X=-y\partial_{x}+x\partial_{y}$ is a vector field for which $f$ is $X$-invariant. With $h^{1}=x^{2}+y^{2}$ and $h^{2}=z$, the functions $h^{1}$ an $h^{2}$ are $X$-invariant. We then construct the data set $\tilde{\mathcal{D}}=\{(h^{1}(x_{i}),h^{2}(x_{i})\}_{i=1}^{r}$. Any machine learning task performed in this feature space will be, by definition, invariant to the flow of $X$, which flow defines the symmetry group of the original data set $\mathcal{D}$.
Another way in which invariant machine learning models can be constructed is by imposing the condition $X(f)=0$ directly in the training of the model $f$. This is discussed in more detail in section 4.4.
By estimating vector fields which satisfy $X(f)=0$, we find seemingly unwanted freedom. This is due to the fact that, assuming $X(f)=0$, $hX(f)=0$ for a nonzero function $h$: the question, then, is which of the two symmetries–the flow of $X$ or the flow of $hX$–is “correct.” By definition, both are symmetries, so that the function $f$ is symmetric with respect to an apparently large class of symmetries. However, a function is $X$-invariant if and only if it is also invariant with respect to $hX$, so that a set of $X$-invariant functions is also a set of invariant functions of $hX$. Thus, where a vector field $X$ can be used to generate $X$-invariant functions, the result is a feature space which is also invariant with respect to $hX$. This point is further expounded upon in section 4.2.2.

### v2

@@ Abstract
Symmetry-informed machine learning can exhibit advantages over machine learning which fails to account for symmetry. In the context of continuous symmetry detection, current state of the art experiments are largely limited to detecting affine transformations. Herein, we outline a computationally efficient framework for discovering infinitesimal generators of multi-parameter group actions which are not generally affine transformations. This framework accommodates the automatic discovery of the number of linearly independent infinitesimal generators. We build upon recent work in continuous symmetry discovery by extending to neural networks and by restricting the symmetry search space to infinitesimal isometries. We also introduce symmetry enforcement of smooth models using vector field regularization, thereby improving model generalization. The notion of vector field similarity is also generalized for non-Euclidean Riemannian metric tensors.
@@
Symmetry Discovery, Machine Learning, Symmetry Enforcement, Killing Vectors, Lie Derivative
{MSCcodes}62-07, 68T01
@@ 1 Introduction
Approaching various machine learning tasks with prior knowledge, commonly in the form of symmetry present in datasets and tasks, has been shown to improve performance and/or computational efficiency [Lyle1, Bergman1, Craven1, tahmasebi2023exact]. While research in symmetry detection in data has enjoyed recent success [LieGAN], the current state of the art leaves room for improvement, particularly with respect to the detection of continuous symmetry [Zhou1]. In this work, we seek to address two issues at hand: primarily, the inability of current methods to detect continuous symmetry beyond affine symmetries; secondly, the apparent increase of difficulty in identifying continuous symmetries when compared with discrete symmetries.
@@
Our methodology of symmetry discovery is summarized as follows. First, given a dataset, certain quantities of interest, hereafter referred to as machine learning functions, are identified, such as an underlying probability distribution which generates the dataset, a classification or regression function, or a dimensionality reduction function. Second, we estimate tangent vector fields which annihilate the machine learning functions. The estimated vector fields themselves are generators of continuous symmetries. Thus, estimating the vector fields provides an indirect estimate of the continuous symmetries.
Vector fields have been used to describe symmetry in the context of machine learning before [Lconv]. However, our use of vector fields differs from previous approaches for two reasons. First, we fully exploit the vector fields as objects which operate on machine learning functions directly. Second, we estimate vector fields associated with continuous symmetries beyond affine transformations.
We also extend the previous work on this subject in nontrivial ways [Shaw1]. While the previous work introduced the notion of symmetry enforcement via the estimation of functions which are invariant with respect to a fixed collection of vector fields, we explore the method of regularization for symmetry enforcement. This method is critical for tasks for which “strict” invariance is not appropriate, or for tasks for which a suitable collection of invariant functions cannot be identified or estimated. A more mathematically rigorous treatment of the estimation of invariant functions is also provided. We also provide a brief overview of infinitesimal generators for multi-parameter group actions in order to highlight that our methodology is suited for the multi-parameter case. We provide more detail and experiments regarding the restriction to special types of vector fields such as Killing vectors: this restriction, as noted in the previous work, can have a particularly positive impact on the estimation of vector fields whose flows are not affine transformations. We generalize the notion of comparing discovered/estimated vector fields to ground-truth vector fields and propose a means to obtain “symmetry scores” to assess the extent to which models are invariant with respect to a vector field. We provide a discussion on the robustness of our symmetry estimation method, and we present all-new experimental results.
@@
This paper is organized as follows. In the next section, we briefly discuss previous work in symmetry detection and enforcement. In Section 3, we discuss vector fields, flows, infinitesimal generators of multi-parameter group actions, and isometries, the mathematical principles that are foundational to our methods. In Section 5, we describe our methods, and in Section 6, we present our experimental results. We then offer concluding remarks and outline future work in the area.
@@ 2.1 Symmetry Discovery
Early work on symmetry detection in machine learning focused primarily on detecting symmetry in image and video data [RaoRuderman, Dickstein1], where symmetries described by straight-line and rotational transformations were discovered. Other work has made strides in symmetry discovery by restricting the types of symmetries being sought. In one case, detection was limited to compact Abelian Lie groups [Cohen1], since such groups exhibit mathematically convenient properties, and used for the purpose of learning disentangled representations. Another case uses meta-learning to discover any finite symmetry group [Zhou1]. Finite groups have also been used in symmetry discovery in representation learning [Anselmi1]. In physics-based applications of machine learning, a method has been constructed which is able to discover any classical Lie group symmetry [Forestano1]. Symmetry discovery of shapes has also been explored [Je1].
@@
Other work has focused on detecting affine transformation symmetries and encoding the discovered symmetries automatically into a model architecture. Three such methods identify Lie algebra generators to describe the symmetries, as we do herein. For example, augerino [Augerino] attempts to learn a distribution over augmentations, subsequently training a model with augmented data. The Lie algebra convolutional network [Lconv], which generalizes Convolutional Neural Networks in the presence of affine symmetries, uses infinitesimal generators represented as vector fields to describe the symmetry. SymmetryGAN [Desai1] has also been used to detect rotational symmetry [LieGAN].
@@
Another notable contribution to efforts to detect symmetries of data is LieGAN. LieGAN is a generative-adversarial network intended to return infinitesimal generators of the continuous symmetry group of a given dataset [LieGAN]. LieGAN has been shown to detect continuous affine symmetries, including transformations from the Lorentz group. It has also been shown to identify discrete symmetries such as rotations by a fixed angle.
While most continuous symmetry detection methods attempt to discover symmetries which are affine transformations, the representation of infinitesimal generators using vector fields has led to the discovery of continuous symmetries which are not affine [Ko1, Shaw1]. We note that our work here is an extension of [Shaw1]. In [Ko1], the domains of image data and partial differential equations are examined in particular, and the symmetries are expressed as flows of vector field infinitesimal generators.However, the approach taken therein differs fundamentally from our approach. In both approaches, the infinitesimal generators are iteratively updated; however, in [Ko1], training data is explicitly transformed by the flow of the vector field, which flow is estimated using a neural network ODE solver. In our work, the infinitesimal generators are taken to act directly upon functions of interest, eliminating the need for data augmentation. Furthermore, the performance of our approach does not depend on the selection of an ODE solver, in contrast to [Ko1].
@@
Continuous symmetry detection differs from discrete symmetry detection [Zhou1] since the condition $f\circ S=f$ must hold for all values of the continuous parameter of $S$. This is corroborated by the increasingly complex methods used to calculate even simple symmetries such as planar rotations [Augerino, Lconv, LieGAN]. Some methods introduce discretization, where multiple parameter values are chosen and evaluated. LieGAN does this by generating various transformations from the same infinitesimal generator [LieGAN]. However, introducing discretization increases the complexity of continuous symmetry detection. A vector field approach addresses the issue of discretization. Our vector field-based method reduces the required model complexity of continuous symmetry detection while offering means to detect symmetries beyond affine transformations.
@@ 2.2 Symmetry Enforcement
We next review methods for enforcing symmetries in machine learning models. Some methods seek to enforce symmetry by augmenting the training dataset according to known symmetries [Bergman1]. This is common practice in image classification tasks, where augmentations include rotated copies of images in the training set. Augerino attempts to enforce symmetry using augmented data, although the symmetries are discovered from the data rather than given a priori. Another established method of enforcing symmetry is feature averaging, which, in contrast to data augmentation, guarantees symmetry enforcement [Lyle1]. Other sources construct target symmetry-invariant and -equivariant models using infinitesimal generators [Lconv, LieGAN]. Some previous work addresses specific cases including the special case of compact groups [Bloem-Reddy1] and equivariant CNNs on homegeneous spaces [Cohen2]. Other works prove universal approximation results for invariant architectures [Maron1, Keriven1, Yarotsky1].
Physics-Informed Neural Networks (PINNs) [Raissi1] may be seen as enforcing a type of symmetry. Here, model training is regularized using differential constraints which represent the governing equations for a physical system. Our regularization method of symmetry enforcement adopts a similar approach, though the differential constraints obtained using infinitesimal generators do not generally have the interpretation of defining governing equations for a physical system. Notably, the topic of symmetry enforcement via regularization has been independently proposed, approximately simultaneously with our own efforts [Otto1].
@@ 3 Background
In this section, we give an overview of vector fields and flows and their connection to symmetry. We also discuss Killing vectors and isometries. Throughout, we assume a dataset $\mathcal{D}=\{x_{1},\dots,x_{r}\}$ with $x_{i}\in\mathbb{R}^{n}$. Various machine learning functions of interest may be defined for a given dataset, such as an underlying probability distribution or class probabilities. The object of symmetry detection in data is to discover functions $S:\mathbb{R}^{n}\to\mathbb{R}^{n}$ that preserve a machine learning function of interest: that is, for a machine learning function $f$, $f\circ S=f$. We consider a continuous symmetry of a particular machine learning function to be a transformation $S$ which is continuously parameterized, such as a rotation in a plane by an angle $\theta$, with $\theta$ being the continuous parameter. We deal exclusively with continuous symmetries.
@@ 3.1 Vector Fields and Flows
We now provide some background on vector fields and their associated flows. We refer the reader to literature on the subject for additional information [Lee1]. Suppose that $X$ is a smooth [^1] That is, a vector field with $\mathcal{C}^{\infty}$ coefficient functions. (tangent) vector field on $\mathbb{R}^{n}$:
@@
For example, for $n=2$, if $f(x,y)=xy$ and $X=y\partial_{x}$, then $(f)=y^{2}. XX$ is also a derivation on the set of smooth functions on $\mathbb{R}^{n}$: that is, as a mapping from $C^{\infty}(\mathbb{R}^{n})$ to itself, it behaves like a derivative in the sense that it is linear and satisfies the product rule. A flow on $\mathbb{R}^{n}$ is a smooth function $\Psi:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ which satisfies
@@
for all $s,t\in\mathbb{R}$ and for all $p\in\mathbb{R}^{n}$. A flow is a 1-parameter group action. An example of a flow $\Psi:\mathbb{R}\times\mathbb{R}^{2}\to\mathbb{R}^{2}$ is
@@
For a given flow $\Psi$, one may define a (unique) vector field $X$ as given in Equation 2, where each function $\alpha^{i}$ is defined as $\alpha^{i}=\left(\frac{\partial\Psi}{\partial t}\right)\bigg|_{t=0}.$ Such a vector field is called the infinitesimal generator of the flow $\Psi$. For example, the infinitesimal generator of the flow given in Equation 4 is $y\partial_{x}+x\partial_{y}$.
@@
Suppose that a solution $\mathbf{x}(t)$ to Equation 5 exists for all $t\in\mathbb{R}$ and for all initial conditions $\mathbf{x}_{0}\in\mathbb{R}^{n}$. Then the function $\Psi:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ given by $\Psi(t,\mathbf{x}_{0})=\mathbf{x}(t)$ is a flow. The infinitesimal generator corresponding to $\Psi$ is $X$. For example, to calculate the flow of $y\partial_{x}+x\partial_{y}$, we solve
@@ 3.2 Vector Fields and Symmetry
We can now connect vector fields and flows with symmetry. A smooth function $f:\mathbb{R}^{n}\to\mathbb{R}$ is said to be $X$-invariant if $X(f)=0$ identically for a smooth vector field $X$. The function $f$ is $\Psi$-invariant if, for all $t\in\mathbb{R}$, $f=f(\Psi(t,\cdot))$ for a flow $\Psi$. If $X$ is the infinitesimal generator of $\Psi$, $f$ is $\Psi$-invariant if and only if $f$ is $X$-invariant. If the function $f$ is a machine learning function for a given data set, our strategy is to identify vector fields $X_{i}$ for which $f$ is invariant. Each vector field $X_{i}$ for which $f$ is invariant is associated, explicitly or implicitly, with a flow $\Psi_{i}$, each of which is a 1-parameter group action, the collection of which generate the multi-parameter group action under which $f$ is invariant. By identifying continuous symmetries by approximately solving $X(f)=0$ for parametric $X$, we boil continuous symmetry discovery down to a tractable optimization problem.
The price of this indirect characterization of continuous symmetry is that the group action is only implicitly represented by its infinitesimal generators. The problem of building machine learning models that are invariant to symmetries corresponding to given infinitesimal generators is called symmetry enforcement, and we describe our results on this in Section 5.6.
3.3 Infinitesimal Generators of Multi-Parameter Group Actions
We emphasize that our methods also apply to symmetries given by multi-parameter group actions, given recent work seeking to compare our method [hu1].
Let $G\subseteq\mathbb{R}^{s}$ be a group, and suppose $G$ acts on $\mathbb{R}^{n}$: that is, for $g_{1},g_{2}\in G$ and for $x\in\mathbb{R}^{n}$, there is a function $\Psi:G\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that (assuming the group operation is vector addition)
(7) $\Psi(\mathbf{0},x)=x,\qquad\Psi(g_{2},\Psi(g_{1},x))=\Psi(g_{1}+g_{2},x).$
The use of the symbol $\Psi$ to denote a multi-parameter group action is not accident, as a flow is a 1-parameter group action. Let $\{v_{i}\}_{i=1}^{s}$ be a basis for the tangent space of $G$ at $\mathbf{0}$, the group identity element. Lastly, let $\sigma$ be a curve in $G$ for which $\sigma(t_{0})=\mathbf{0}$ and $\dot{\sigma}(t_{0})=v_{i}$ for $t_{0}\in\mathbb{R}$. The infinitesimal generator $X_{i}$ corresponding to $v_{i}$ is given by
(8) $X_{i}=\left(\dfrac{d}{dt}\Psi(\sigma(t),x)\right)\bigg|_{t=t_{0}}.$
For example, consider the group $G=\mathbb{R}^{3}$ acting on $\mathbb{R}^{2}$ via
$\Psi((a,b,\theta),(x,y))=\left(x\cos(\theta)-y\sin(\theta)+a,x\sin(\theta)+y\cos(\theta)+b\right).$
Given the following three curves, $\sigma_{a}(t)=(t,0,0),\qquad\sigma_{b}(t)=(0,t,0),\qquad\sigma_{\theta}(t)=(0,0,t),$ we find that $X_{a}=\dfrac{d}{dt}\left(t,0\right)|_{t=0}=\partial_{x},\qquad X_{b}=\dfrac{d}{dt}\left(0,t\right)|_{t=0}=\partial_{y},$
$X_{\theta}=\dfrac{d}{dt}\left(x\cos(\theta)-y\sin(\theta),x\sin(\theta)+y\cos(\theta)\right)|_{t=0}=-y\partial_{x}+x\partial_{y}.$ For each of these vector fields, a corresponding flow can be computed, which flows we call $\Psi_{a}$, $\Psi_{b}$, and $\Psi_{\theta}$, respectively. In terms of the original parameters, these flows are given as $\Psi_{a}(a,(x,y))=(x+a,y),\Psi_{b}(b,(x,y))=(x,y+b),$ and $\Psi_{\theta}=\left(x\cos(\theta)-y\sin(\theta),x\sin(\theta)+y\cos(\theta)\right)$. While each of these flows are, individually, 1-parameter group actions, it is clear that the infinitesimal generators $X_{a}$, $X_{b}$, and $X_{\theta}$ are the infinitesimal generators for the multi-parameter group action given in Equation (7). Thus, discovering vector field infinitesimal generators which annihilate a fixed (smooth) function applies to multi-parameter group actions and not solely to 1-parameter group actions.